OpenStudy (samsan9):

Help please! Solve each system. In your work, identify the shape of the graph of each equation. x^2+y^2+2x+2y=0 x^2+y^2+4x+6y+12=0

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ x^2+y^2+2x+2y=0 }\) \(\normalsize\color{blue}{ x^2+2x+y^2+2y=0 }\) \(\normalsize\color{blue}{ x^2+2x\color{red}{+1 } +y^2+2y\color{green}{+1 } =0 \color{red}{+1 } \color{green}{+1 } }\) \(\normalsize\color{blue}{ x^2+2x+1+~y^2+2y+1=2 }\) \(\normalsize\color{blue}{ \underline{x^2+2x+1}+ \underline{y^2+2y+1}=2 }\) \(\normalsize\color{blue}{ (x+1)^2+(y+1)^2=2 }\) \(\normalsize\color{blue}{ (x+1)^2+(y+1)^2=\sqrt{2}^{2} }\)

OpenStudy (solomonzelman):

Compare this one to the equation of the circle, \(\normalsize\color{blue}{ (x-h)^2+(y-k)^2=r ^{2} }\) where `(h,k)` is the center.

OpenStudy (solomonzelman):

The other one, \(\normalsize\color{blue}{ x^2+y^2+4x+6y+12=0 }\) \(\normalsize\color{blue}{ x^2+4x+12+y^2+6y=0 }\) \(\normalsize\color{blue}{ x^2+4x+4+8+y^2+6y=0 }\) \(\normalsize\color{blue}{ x^2+4x+4+y^2+6y+8=0 }\) \(\normalsize\color{blue}{ x^2+4x+4+y^2+6y+8\color{red}{+1}=0\color{red}{+1} }\) \(\normalsize\color{blue}{ \underline{ x^2+4x+4 } +\underline{ y^2+6y+9}=1 }\) \(\normalsize\color{blue}{ (x+2)^2 +(y+3)^2=1 }\) \(\normalsize\color{blue}{ (x+2)^2 +(y+3)^2=1^2 }\)

OpenStudy (solomonzelman):

Compare this one also to \(\normalsize\color{blue}{ (x-h)^2 +(y-k)^2=r^2 }\) . Good LUCK !

OpenStudy (samsan9):

thank you so much

OpenStudy (solomonzelman):

Do you see the centers and the radiuses ?

OpenStudy (samsan9):

um sure can you still post them my parents are making me leave the house for a few minutes sadly but i would appreciate it

OpenStudy (solomonzelman):

Yes :) radius center Circle 1 ─ √2 (-1,-1) Circle 2 ─ 1 (-2,-3)

OpenStudy (solomonzelman):

Good luck !

OpenStudy (solomonzelman):

ntw I am also about to leave:) Basketball

OpenStudy (solomonzelman):

btw (not ntw)

OpenStudy (samsan9):

ok good luck and thanks :D